vincent is 10 years older than john paul. 5 years ago, his age was twice that of john paul's age. findd their present ages.
Answer by mananth(16946)

vincent is 10 years older than john paul. 5 years ago, his age was twice that of john paul's age. findd their present ages.
paul x years
Vincent = x+10 years
..
5 years ago
paul x-5 years
Vincent = x+10-5=x+5
..
x+5=2(x-5)
x+5 = 2x-10
2x-x=10+5
x= 15 paul's age
...
Vincent's age =10+15 = 25 years
...
m.ananth@hotmail.ca